# Molecular Identification and Peak Assignment: Leveraging Multi-Level Multimodal Alignment on NMR

### v1

@@ Abstract
Nuclear magnetic resonance (NMR) spectroscopy plays an essential role across various scientific disciplines, providing valuable insights into molecular dynamics and interactions. Despite the promise of AI-enhanced NMR prediction models, challenges persist in the interpretation of spectra for tasks such as molecular retrieval, isomer recognition, and peak assignment. In response, this paper introduces Multi-Level Multimodal Alignment with Knowledge-Guided Instance-Wise Discrimination (K-M3AID) to establish meaningful correspondences between two heterogeneous modalities: molecular graphs (structures) and NMR spectra. In particular, K-M3AID employs a dual-coordinated contrastive learning architecture, and incorporates a graph-level alignment module, a node-level alignment module, and a communication channel. Notably, the framework introduces knowledge-guided instance-wise discrimination into contrastive learning within the node-level alignment module, significantly enhancing accuracy in cross-modal alignment. Additionally, K-M3AID showcases its capability of meta-learning by demonstrating that skills acquired during node-level alignment positively impact graph-level alignment. Empirical validation underscores K-M3AID’s effectiveness in addressing multiple zero-shot tasks, offering a promising solution to bridge the gap between structural information and spectral data in complex NMR scenarios.
@@ 1 Introduction
Nuclear magnetic resonance (NMR) spectroscopy has found a board applications in various scientific domains, such as organic chemistry, pharmaceuticals, biochemistry, food science, and material science, by providing insights into molecular dynamics and interactions (Gunther & Gunther (1994); Claridge (2016); Yu et al. (2021)). The positions and shapes of signal peaks in NMR spectra, influenced by through-bond and through-space interactions, serve as ”fingerprints” for deducing atomic connectivity, relative stereochemistry, and conformations. The conventional approach for elucidating molecular structures and attributing peaks has long relied on manual determination by organic chemists (Guan et al. (2021)). However, the interpretation of NMR spectra is not straightforward, particularly when dealing with isomers and complex molecules consisting of multiple stereogentic (chiral) centers (Wu et al. (2023); Chhetri et al. (2018)). Even an expert chemist may encounter significant difficulties in accurately assigning isomeric compounds with extremely similar NMR spectra due to this complexity (Nicolaou & Snyder (2005)).
Recent advancements in AI-enhanced NMR prediction models show promise in generating spectra from candidate structures (Chen et al. (2020); Jonas et al. (2022); Kuhn (2022)). However, real-world scenarios often demand spectral interpretation before detailed structural information is available. Ideally, the interpretation of spectra should transition from spectral data to structural elucidation. Consequently, these prediction models are not applicable in highly demanding real-world tasks, such as molecular retrieval, candidate ranking, and peak assignment (see Figure 1.a). In scenarios with limited knowledge of an unknown compound, such as identifying a compound from a plant, molecular retrieval is crucial, as the retrieved analogs can aid in structural hypothesis formulation. In cases where pre-existing knowledge is available, such as identifying reaction products, the quantitative ranking of candidate possibilities becomes essential. Peak assignment represents the decisive stage in the structural identification process, providing detailed insights into the local atomic context. These prediction models fall short in identifying compounds without prior knowledge, quantitatively assessing candidates (especially in isomer recognition), and guiding accurate peak assignments. How might one employ or interrogate these advanced AI methodologies to enhance molecular retrieval, candidate ranking, and peak assignment in NMR spectra interpretation?
NMR spectroscopy can be viewed as a modality, capturing molecular structural and electronic details through an NMR spectrometer and expressing this information in sequence format (chemical shifts). On the other hand, these structural and electronic information can also be presented in a molecular graph, which encapsulates molecular attributes in a graph-based format. The tasks of determining molecular structure and assigning peaks require clear correspondence across these two heterogeneous modalities: NMR spectra and molecular graphs. This imperative is met by multimodal learning, particularly multimodal alignment (see fundamentals of multimodal alignment in Preliminary  2). However, multimodal alignment commonly encounters challenges in addressing data heterogeneity, semantic gaps among modalities, collecting paired data, designing effective cross-modal attention mechanisms, and formulating appropriate evaluation metrics (Liang et al. (2023); Jabeen et al. (2023)).
Molecules is composed of atoms, and each molecule, with its unique arrangement of atoms and bonding patterns, gives rise to a distinct spectrum. Consequently, it is impractical to incorporate all molecules and their spectra in the training dataset. One of the potential solutions is to take into account the constrained possibilities of atomic diversity for interpreting spectra. This, in turn, motivates us to leverage the power of AI for interpreting spectra at the atomic level, establishing a solid foundation for molecular identification. This approach aligns tightly with the core principles of meta-learning (see principles of meta-learning in Preliminary  2). However, meta-learning typically concerns in meta-overfitting, distance evaluation metrics, model interpretablity, transferability and generalization (Huisman et al. (2021); Feng et al. (2022); Ma et al. (2022); Vettoruzzo et al. (2023)).
In view of these challenges and opportunities, we propose a novel framework K-M3AID (Multi-Level Multimodal Alignment with Knowledge-Guided Instance-Wise Discrimination), aiming at achieving reliable analog retrieval, candidate ranking, and peak assignment in the interpretation of NMR spectra (see Figure  1.b). The overview of our K-M3AID framework is a dual-coordinated contrastive learning architecture, which contains three key components: graph-level alignment module, node-level alignment module, and communication channel. Graph-level alignment module establishes the correspondences of molecules with their individual ${}^{13}$ NMR spectra. Given that each unique molecule gives rise to a distinct spectral signature, this module is equipped with a straightforward cross-entropy loss designed for effective contrastive learning. Node-level alignment module aligns each C atom within the molecules with their signal peaks on the spectrum. In contrast to the diverse and distinctive molecular spectral signatures, many atoms exhibit chemical symmetric and magnetic equivalence within the same molecule, corresponding to the same peaks. Meanwhile, atoms with different local surroundings can still presents significant similarity on the spectrum, which introduces heightened level of complexity. In view of these complex scenarios, we come up with knowledge-guided instance wise discrimination based contrastive learning in graph-level alignment module (see Figure 4). The communication channel dynamically adjusts the flow of gradients between the node encoder and the graph encoder from two modules during the training process.
In summary, our contribution comprises three major aspects: Conceptually: we integrate two architectural levels alignment, graph- and node-levels, into the K-M3AID framework. This integration facilitates rapid adaptation, significantly boosting the efficiency of learning for multimodal zero-shot tasks. Methodologically: We present knowledge-guided instance wise discrimination for cross-modal contrastive learning, which take advantage of continuous and domain-specific features with natural ordering. To the best of our knowledge, this is the first to demonstrate knowledge-guided instance wise discrimination based cross-modal contrastive learning, which transforms a discrete comparison into a continuous one. Empirically: We demonstrate the effectiveness of K-M3AID through its successful application to multiple zero-shot tasks, including molecular retrieval, isomer recognition, and peak assignment.
@@ 2 Preliminaries
Multimodal Alignment: Multimodal alignment, as defined in the literature (Baltrusaitis et al. (2017)), involves establishing relationships and correspondences among sub-components of instances from two or more modalities. A typical example is identifying specific regions in an image that correspond to words or phrases in a given caption (Karpathy & Fei-Fei (2015)). This approach offers numerous benefits, including enhanced data interpretation, heightened accuracy and robustness, overcoming limitations of single-modal systems, and better addressing real-world complexity (Baltrusaitis et al. (2017), Summaira et al. (2021), Akkus et al. (2023)). CLIP (Contrastive Language-Image Pretraining) (Radford et al. (2021); Li et al. (2021)) is one of the most widely adopted frameworks for multimodal alignment. As highlighted in the introduction, molecular information originates from diverse sources such as molecule graphs and NMR spectroscopy. To leverage effective alignments of this multifaceted information across different modalities, we adopt CLIP framework with graph neural networks (GNN) (Xu et al. (2018), Wu et al. (2022)) to encoder molecular information and neural network encoders (Serra et al. (2018)) to encapsulate NMR information.
@@
Meta-Learning: Meta-learning is defined as learning how to learn across tasks (Vilalta & Drissi (2002)). More specifically, it leverages skills previously acquired from related tasks to the current one (Lake et al. (2017). With more skills learned, learning new skills becomes easier, requiring fewer examples and less trial-and-error (Vanschoren (2018); Finn et al. (2017)). A meta-learner is trained on a diverse set of object recognition tasks. During this training, it learns common features, patterns, and strategies for recognizing objects. Once trained, when presented with a new, previously unseen object category, the meta-learner can rapidly adapt and achieve high recognition accuracy, leveraging the knowledge it acquired from the diverse training tasks to perform in this novel recognition task (Finn et al. (2017). A profound understanding of atom properties allows us to extend our vision to previously unseen molecules, aligning with the principles of meta-learning in artificial intelligence.
@@
Contrastive Learning: Contrastive learning focus on learning how to distinguish between similar and dissimilar items (Le-Khac et al. (2020b); Jaiswal et al. (2021); Liu et al. (2021)). A fundamental aspect of this process involves instance-wise discrimination (Wu et al. (2018)). Models incorporating instance-wise discrimination not only foster an understanding of the inherent data structure but also enhance generalization capabilities. This is attributed to the contrastive learning approach, which prioritizes distinguishing between instances rather than memorizing specific labeled examples. In chemistry, each molecule/atom is treated as a distinct instance. The learning algorithm focuses on distinguishing each molecule/atom based on its context.
@@ 3 Experiments
To thoroughly evaluate the performance of K-M3AID, we compare it with other baselines through various zero-shot downstream tasks, such as molecular retrieval, isomer recognition and peak assignment.
3.1 Experimental Setting
3.1.1 Pre-training
Dataset: We use ${}^{13}$ NMR spectra of about 20,000 molecules sourced from nmrshiftdb2 (Steinbeck et al. (2003)), a public access database that contains NMR spectra of organic molecules. In the collected dataset, molecule are aligned with their respective ${}^{13}$ NMR spectra, and atomic alignments with peaks are also included. Notably, the dataset contains 12,771 molecules with fewer than 10 carbon atoms, 7,043 molecules featuring carbon atom counts ranging from 10 to 20, and 1,138 molecules incorporating more than 20 carbon atoms. The quality of the dataset was further validated by experienced organic chemists. We randomly sample 80% of the molecules for training and the rest for evaluation.
3.1.2 Zero-Shot Molecular Retrieval
Dataset: We randomly collected 1 million molecules from PubChem (Kim et al. (2023)) to form a molecular reference library. Subsequently, we carefully selected 1000 spectra, ensuring that they had not appeared in the training dataset, from an external dataset to serve as query spectra. Following this, the corresponding molecules associated with these 1000 spectra were added into the existing reference library.
Evaluation: We perform molecular retrieval using each of the selected spectra to determine if the correct corresponding molecular entity can be retrieved from the reference library. The model’s performance is assessed at top-1% accuracy, as well as at accuracy of 5%, 10%, and 25%.
3.1.3 Zero-Shot Isomer Recognition
Dataset: We categorize isomers from the validation dataset to guarantee their absence from the training dataset. To assess effectiveness, we perform isomer recognition on each isomer group containing at least 10 molecules. Within the same group, isomers may be structural or spatial isomers of each other. Structural isomers refer to molecules with the same molecular formula but different structural arrangements of atoms, resulting in distinct chemical structures. On the other hand, spatial isomers, also known as stereoisomers, have the same molecular formula and arrangement of atoms but differ in the spatial orientation of their atoms in three-dimensional space, leading to different stereoisomeric forms. (For an elucidation of an isomer group and in-depth insights into isomers with NMR, please refer to the details provided in Appendix D.)
3.1.4 Zero-Shot Peak Assignment
3.1.5 Baselines

### v2

@@ Abstract
Nuclear magnetic resonance (NMR) spectroscopy plays an essential role in deciphering molecular structure and dynamic behaviors. While AI-enhanced NMR prediction models hold promise, challenges still persist in tasks such as molecular retrieval, isomer recognition, and peak assignment. In response, this paper introduces a novel solution, Multi-Level Multimodal Alignment with Knowledge-Guided Instance-Wise Discrimination (K-M3AID), which establishes correspondences between two heterogeneous modalities: molecular graphs and NMR spectra. K-M3AID employs a dual-coordinated contrastive learning architecture with three key modules: a graph-level alignment module, a node-level alignment module, and a communication channel. Notably, K-M3AID introduces knowledge-guided instance-wise discrimination into contrastive learning within the node-level alignment module. In addition, K-M3AID demonstrates that skills acquired during node-level alignment have a positive impact on graph-level alignment, acknowledging meta-learning as an inherent property. Empirical validation underscores K-M3AID’s effectiveness in multiple zero-shot tasks.
Machine Learning, ICML
@@ 1 Introduction
Nuclear magnetic resonance (NMR) spectroscopy has found broad applications in various scientific domains, such as chemistry, environmental science, food science, material science, and pharmaceuticals, by providing insights into molecular dynamics and structures (Gunther & Gunther, 1994; Claridge, 2016; Yu et al., 2021). The details of NMR spectra can be influenced by through-bond and through-space interactions, serving as ”fingerprints” to deduce atomic connectivity, relative stereochemistry, and conformations. The conventional approach for elucidating molecular structures and attributing peaks has long relied on manual determination by organic chemists (Guan et al., 2021). However, the interpretation of NMR spectra is not straightforward, particularly when dealing with isomers and complex molecules consisting of multiple stereogenic (chiral) centers (Wu et al., 2023; Chhetri et al., 2018). Even an expert chemist may encounter significant difficulties in accurately assigning isomeric compounds with extremely similar NMR spectra due to this complexity (Nicolaou & Snyder, 2005).
@@
While recent AI-enhanced NMR spectrum prediction models show promise in generating spectra from candidate structures (Chen et al., 2020; Jonas et al., 2022; Kuhn, 2022), these models still face challenges in peak assignment due to their high error tolerance and a lack of precise point-to-point guidance. Since peak assignment is a determining step in isomer recognition, these models fall short in achieving accurate isomer recognition. Another contributing factor is the absence of quantitative ranking for candidate isomers in their implementation. In addition, the success of these models requires a good level of prior knowledge of molecular structures to construct candidates. However, real-world practice often demands spectral interpretation before detailed structural information is available. For instance, when identifying an unknown compound from a plant, there is limited or no knowledge of this compound. Thus, the interpretation of spectra should transition from spectral data to structural elucidation. How might one employ a series of advanced AI methodologies to simplify NMR spectra interpretation, particularly in tasks like molecular retrieval, candidate ranking, and peak assignment (see Figure  1.a)?
Multimodal alignment presents an opportunity to address these challenges in the context of two heterogeneous modalities: NMR spectrum and molecular graphs. The NMR spectrum is a sequence-based chemical modality that captures molecular structural and electronic details through an NMR spectrometer, translating such information into NMR peaks. A molecular graph can encapsulate molecular structural and electronic information through the arrangement of nodes and edges, along with their respective attributes. The analysis of molecular structure and peak assignment requires clear correspondence across these two heterogeneous modalities. This imperative is met by the principles of multimodal alignment (see Preliminary  2). However, multimodal alignment commonly encounters challenges in addressing data heterogeneity, semantic gaps among modalities, collecting paired data, designing effective cross-modal attention mechanisms, and formulating appropriate evaluation metrics (Liang et al., 2023; Jabeen et al., 2023).
Meta-learning can serve as an accelerating approach to facilitate the learning process in the context of the multi-view nature of molecules. Molecules are distinguished by the distinctive configuration of atoms coupled with bonding patterns, giving rise to distinct spectra. As molecular diversity is extensive, it is impractical to include all molecules and their spectra in a training dataset. However, the corresponding atomic diversity is comparatively constrained. Thus, it is a sound approach to analyze molecular structures by interpreting spectra at the atomic level. This approach aligns tightly with the core principles of meta-learning (see Preliminary  2). However, meta-learning typically concerns meta-overfitting, distance evaluation metrics, model interpretability, transferability, and generalization (Huisman et al., 2021; Feng et al., 2022; Ma et al., 2022; Vettoruzzo et al., 2023).
In light of these challenges and opportunities, we propose a novel framework, K-M3AID (Multi-Level Multimodal Alignment with Knowledge-Guided Instance-Wise Discrimination), aiming to achieve reliable analog retrieval, candidate ranking, and peak assignment in the interpretation of NMR spectra (see Figure  1.b). The overview of our K-M3AID framework features a dual-coordinated contrastive learning architecture, comprising three key components: a graph-level alignment module, a node-level alignment module, and a communication channel. The graph-level alignment module establishes correspondences between molecules and their individual ${}^{13}$ NMR spectra. Given that each unique molecule produces a distinct spectral signature, this module employs a straightforward cross-entropy loss for effective contrastive learning. The node-level alignment module aligns each Carbon atom within the molecules with their signal peaks on the spectrum. Unlike the diverse and distinctive molecular spectral signatures, many atoms exhibit chemical symmetry and magnetic equivalence within the same molecule, corresponding to the same peaks. However, atoms with different local surroundings can still present significant similarity on the spectrum, introducing a heightened level of complexity. To address these complex scenarios, we introduce knowledge-guided instance-wise discrimination based on contrastive learning in the graph-level alignment module (see Figure 2). The communication channel dynamically adjusts the flow of gradients between the node encoder and the graph encoder from two modules during the training process.
In summary, our contribution encompasses three significant aspects: Conceptually: We integrate cross-modal alignment at two architectural levels, namely graph and node levels, within the K-M3AID framework. This integration facilitates rapid adaptation, significantly boosting the efficiency of learning for zero-shot tasks. Methodologically: We introduce knowledge-guided instance-wise discrimination for cross-modal contrastive learning, leveraging continuous and domain-specific features with inherent natural ordering. To the best of our knowledge, this is the first demonstration of knowledge-guided instance-wise discrimination-based cross-modal contrastive learning, transforming discrete comparisons into a continuous paradigm. Empirically: We substantiate the effectiveness of K-M3AID through its successful application to various zero-shot tasks, including molecular retrieval, isomer recognition, and peak assignment.
@@ 2 Preliminaries
Multimodal Alignment: Multimodal alignment, as defined in the literature (Baltrusaitis et al., 2017), involves establishing relationships and correspondences among sub-components of instances from two or more modalities. A typical example is identifying specific regions in an image that correspond to words or phrases in a given caption (Karpathy & Fei-Fei, 2015). This approach offers numerous benefits, including enhanced data interpretation, heightened accuracy and robustness, overcoming limitations of single-modal systems, and better addressing real-world complexity (Baltrusaitis et al., 2017), (Summaira et al., 2021), (Akkus et al., 2023). CLIP (Contrastive Language-Image Pretraining) (Radford et al., 2021; Li et al., 2021) is one of the most widely adopted frameworks for multimodal alignment. As highlighted in the introduction, molecular information originates from diverse sources such as molecule graphs and NMR spectroscopy. To leverage effective alignments of this multifaceted information across different modalities, we adopt the CLIP framework with graph neural networks (GNN) (Xu et al., 2018), (Wu et al., 2022) to encode molecular information and neural network encoders (Serra et al., 2018) to encapsulate NMR information.
@@
Meta-Learning: Meta-learning is defined as the process of learning how to learn across tasks (Vilalta & Drissi, 2002). More specifically, it leverages skills previously acquired from related tasks to the current one (Lake et al., 2017). With more skills learned, acquiring new ones becomes easier, requiring fewer examples and less trial-and-error (Vanschoren, 2018; Finn et al., 2017). A meta-learner is trained on a diverse set of object recognition tasks. During this training, it learns common features, patterns, and strategies for recognizing objects. Once trained, when presented with a new, previously unseen object category, the meta-learner can rapidly adapt and achieve high recognition accuracy, leveraging the knowledge acquired from the diverse training tasks to perform in this novel recognition task (Finn et al., 2017). A profound understanding of atom properties allows us to extend our vision to previously unseen molecules, aligning with the principles of meta-learning in artificial intelligence.
@@
Contrastive Learning: Contrastive learning focuses on discerning similarities and differences between items (Le-Khac et al., 2020b; Jaiswal et al., 2021; Liu et al., 2021). A fundamental aspect of this process involves instance-wise discrimination (Wu et al., 2018). Models incorporating instance-wise discrimination not only foster an understanding of the inherent data structure but also enhance generalization capabilities. This is attributed to the contrastive learning approach, which prioritizes distinguishing between instances rather than memorizing specific labeled examples. In chemistry, each molecule/atom is treated as a distinct instance, and the learning algorithm focuses on distinguishing each molecule/atom based on its context.
3 Our Method
In this section, we firstly introduce Knowledge-Guided Instance-Wise Discrimination. Then, we present the architecture of the K-M3AID framework, an end-to-end system designed for multi-level multimodal alignment, along with its loss function.
Figure 2: Knowledge-Guided Instance-Wise Discrimination Mechanism. $K_{i}$ and $K_{j}$ represents the corresponding knowledge span label for $i^{th}$ and $j^{th}$ items.
Knowledge Span, which we define as a continuous and domain-specific feature exhibiting natural ordering and offering guidance, can potentially offer valuable insights into the contrastive learning process. As such, we propose a novel and general approach to contrastive learning called Knowledge-Guided Instance-Wise Discrimination (see Figure  2). This approach expands the scope of contrastive learning from confined comparisons (pre-determined negative and positive pairs) to unrestricted comparisons (no need for pre-determination). This extension removes the necessity of explicitly defining such pairs, thus mitigating the potential introduction of human bias.
Suppose $\mathcal{M}$ is the set of instances. $\mathcal{A}\subset\mathbb{R}^{d_{1}}$ is the set of tunable instances’ embeddings in modality A, $\mathcal{B}\subset\mathbb{R}^{d_{1}}$ is the set of tunable instances’ embeddings in modality B, and $\mathcal{K}\subset\mathbb{R}^{d_{2}}$ is the corresponding fixed knowledge span label that can guide the relative distance learning between components in $\mathcal{A}$ and $\mathcal{B}$. Thus, the size of $\mathcal{A}$, $\mathcal{B}$, $\mathcal{K}$ are $|\mathcal{M}|$, respectively.
3.2 Architecture & Contrastive Learning Loss
The K-M3AID framework is a dual-CLIP architecture (see Figure 1), consisting of three critical components: a graph-level alignment module, a node-level alignment module, and a communication channel. The graph-level alignment module adopts a gradient-asymmetric CLIP mechanism. While two unimodal encoders work in conjunction, only the from-scratch graph encoder (GIN, (Xu et al., 2018)) undergoes dynamic training throughout the process; the pre-trained spectrum encoder (Yang et al., 2021) remains fixed. Both encoders are complemented by dedicated projection layers, facilitating the mapping of embeddings into a joint space. The node-level alignment module adopts a gradient-symmetric CLIP mechanism. It is equipped with two from-scratch unimodal encoders, the node encoder and the peak encoder, as well as their dedicated projection layers. The graph encoder in the graph-level alignment module shares part of the weights with the node encoder in the node-level alignment module, serving as the communication channel.
The synergy between these two modules is pivotal, collectively contributing to the loss function, expressed as
$L=CL_{graph}+CL_{node},$ (8)
$=-\text{log}(\text{softmax}(\delta(x_{i},y_{i}))$ (10)
Thus, the total $CL_{graph}$ is expressed as following:
$CL_{graph}=\frac{1}{N}\displaystyle\sum_{1\leq i\leq N}CL_{graph}(i)$ (11)
@@ 4 Experiments
To thoroughly evaluate the performance of K-M3AID, we compare it with other baselines across various zero-shot downstream tasks, including molecular retrieval, isomer recognition, and peak assignment. Please refer to the detailed settings of pre-training and downstream tasks in Appendix B.
4.1 Chosen Knowledge Span-ppm
${}^{13}$ NMR uncovers molecular structures by providing the chemical environments of carbon atoms and their magnetic responses to external fields, quantifying these features in parts per million (ppm) relative to a reference compound like tetramethylsilane (TMS), simplifying comparisons across experiments. Thus, continuous peak positions, measured in ppm, can serve as a robust knowledge span to facilitate instance-wise discrimination for this contrastive learning task.
For the node-level alignment module, $\mathcal{A}$ is the set of node embeddings for Carbon atoms in the molecular graph modality, and $\mathcal{B}$ is the set of peak embeddings for respective Carbon atoms in the NMR modality. $\mathcal{K}$ is the set of ppm values for each corresponding Carbon atom in $\mathcal{A}$ and $\mathcal{B}$. Suppose $ppm_{i}$ is the ppm for the $i^{th}$ Carbon Atom, and $ppm_{j}$ is the corresponding ppm for the $j^{th}$ peak. $d(\cdot,\cdot)$ is then defined as follows:
$d(\mathcal{K}_{i},\mathcal{K}_{j})$ $=d(ppm_{i},ppm_{j})$ (12)
$=softmax(\frac{\tau_{2}}{|ppm_{i}-ppm_{j}|+\tau_{1}})$ (13)
where $\tau_{1}$ and $\tau_{2}$ are temperature hyper-parameter. For further discussion of selection about $\tau_{1}$ and $\tau_{2}$, please refer to Appendix C.2. Then, the final form of contrastive loss for node-level alignment according to Equation 5 and Equation 7 is as following:
$CL_{node}$ $=-\frac{1}{|\mathcal{M}|}\sum_{j=1}^{|\mathcal{M}|}d(\mathcal{K}_{i},\mathcal{K}_{j})\cdot log\frac{e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}{\sum_{k=1}^{|\mathcal{M}|}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{k})}}$ (14)
Here, $i$ and $j$ are indices of atoms. $\mathcal{A}_{i}\in\mathcal{A}$ represents the embedding of $i-th$ atom in modality A while $\mathcal{B}_{i}\in\mathcal{B}$ represents the embedding of $i-th$ atom in modality B, $\mathcal{K}_{j}\in\mathcal{K}$ represents peaks, and $th$ is the abbreviation for the threshold.
4.2 Baselines
No Communication: In contrast to the communicative mechanism of K-M3AID, one of the baselines is established without the utilization of a communication channel (denoted as No Comm.).